OpenStudy (anonymous):

i got 4.5 to both answers so i want to know if i did ok wo congruent equilateral triangles have sides of 23 mm and 4x + 5 mm. Find the value of x. The total number of degrees in a regular pentagon is 540°. If a corresponding angle in a congruent regular pentagon is 16x - 36°, find the value of x.

OpenStudy (anonymous):

hi suzy

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

where are you from?

OpenStudy (anonymous):

are you gonna help me im from the u.s.

OpenStudy (anonymous):

sure i'll help you

OpenStudy (anonymous):

equilateral means?

OpenStudy (anonymous):

equal sides right?

OpenStudy (anonymous):

so 23mm = 4x + 5

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

4x=23-5

OpenStudy (anonymous):

23=4x+5 18=4x x=4.5 my work

OpenStudy (anonymous):

4x=18 and x=4.5(final answer)

OpenStudy (anonymous):

and the second one

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

how old are you btw? what grade question is this coming from?

OpenStudy (anonymous):

i did these sorts of question in year7 lol

OpenStudy (anonymous):

im 21 r u from china?

OpenStudy (anonymous):

oh cool i'm 17, i live in hong kong

OpenStudy (anonymous):

so some people have it differently

OpenStudy (anonymous):

strictly speaking, it's china you can say lol

OpenStudy (anonymous):

yea i guessed right wel thank u

OpenStudy (anonymous):

show me your working for this 2nd one, i'll help you check

OpenStudy (anonymous):

is this for the SAT btw? i'm gonna take it on may 8th :P

OpenStudy (anonymous):

540/5=16x-36 108=16x-36 72=16x x=4.5

OpenStudy (anonymous):

hey looks right

OpenStudy (anonymous):

don't forget your degrees

OpenStudy (anonymous):

ok and no its for a diploma thank you

OpenStudy (anonymous):

4.5degrees

OpenStudy (anonymous):

i'm gonna give you a medal ^_^

OpenStudy (anonymous):

how generous

OpenStudy (anonymous):

if you have any other Qs feel free to message me

OpenStudy (anonymous):

ok thx